# Mpemba effect in a two-dimensional bistable potential

### 25 Mar 2026 v1

@@ IV.1.2 $a_{m}$ for $m=2,3,4,5$
Figure 5(a) shows $\beta_{\mathrm{ini}}$ dependence of $a_{2}$ and Fig. 5(b) shows $a_{3}$, $a_{4}$, and $a_{5}$. It confirms the existence of a peak of $a_{2}$ against $\beta_{\mathrm{ini}}$, which is the necessary condition to observe the Mpemba effect. Remarkably, we expect that the overdamped Langevin system confined in a two-dimensional asymmetric potential exhibits the peak of $a_{2}$ without the introduction of a hard wall. This result contrasts with that in the one-dimensional asymmetric potential, where the Mpemba effect disappears if the system contains no hard walls [53, 54].
@@ IV.2 Higher-order contributions in the spectrum decomposition
Figure 6(a) shows the initial distribution $P_{\mathrm{ini}}$, q. (59), for $\beta_{\mathrm{ini}}=0.82$ with various $N_{\mathrm{tr}}(=3,5,7)$. Although the approximation is not accurate for $N_{\mathrm{tr}}=3$, it becomes sufficiently accurate when $N_{\mathrm{tr}}=7$. Figure 6(b) plots the time evolution of the distribution function for $N_{\mathrm{tr}}=7$. For clarity, we plot the difference from the final state $P_{\mathrm{eq}}(r,\beta)$. Since the smallest eigenvalue is $\tilde{\lambda}_{2}=0.20249779968052614$, the distribution function is found to converge to $P_{\mathrm{eq}}(r,\beta)$ for $t\gg 1/\tilde{\lambda}_{2}\approx 5$. Furthermore, as shown in Appendix C, the Mpemba effect is still observed in terms of the difference in the KL divergence even when higher-order modes are included. Therefore, Eq. (II.3.1) provides a valid description of the dynamics.
@@ IV.3 Condition of the crossing of the KL divergence
Although we have confirmed the existence of a peak of $a_{2}$ against $\beta_{\mathrm{ini}}$, this may not be sufficient to observe the Mpemba effect. To observe the Mpemba effect, we need a crossing condition for the observable. In this subsection, we clarify the sufficient condition to observe the Mpemba effect.
Among many possible observables, we choose the KL divergence defined as
@@
which is a typical monotone measure, where $\delta P:=P(\bm{r},t)-P_{\mathrm{eq}}(\bm{r})$. If we are interested in the late stage of relaxation dynamics or the deviation from the initial state and the final state is small, $D^{\mathrm{KL}}(P(t)||P_{\mathrm{eq}})$ can be approximated as
@@
for two different initial conditions A and B. See Appendix B for the derivation of Eq. (64). This quadratic form is valid only when $P(r,t)$ is sufficiently close to equilibrium; in particular, it does not reproduce the exact value at $t=0$ when the initial condition is another equilibrium distribution. Therefore, the condition Eq. (64) should be regarded as a necessary (but not always sufficient) condition for crossing within the validity of the truncated expansion.
@@
We plot the time evolution of the KL divergence $\Delta D(\beta^{\sharp},\beta^{*}):=D^{\mathrm{KL}}(P(\beta^{\sharp},\beta;t)||P_{\mathrm{eq}}(\beta^{\sharp}))-D^{\mathrm{KL}}(P(\beta^{*},\beta;t)||P_{\mathrm{eq}}(\beta))$ during a cooling process starting from $\beta_{\mathrm{ini}}^{\sharp}=0.70$ and $\beta_{\mathrm{ini}}^{*}=0.82$ to $\beta=1$ in Fig. 7 (see the solid line). It is apparent to observe the crossing around $t\approx 1$, which exhibits the Mpemba effect. We now compare the KL divergence, including higher-order terms, by using the expansion of $P(r,t)$ truncated at $N_{\mathrm{tr}}=7$ as the dashed line. The dotted line represents the semi-analytic result of Eq. (93) (see Appendix C). From these results, the two-mode approximation works well, although the crossing time depends on $N_{\mathrm{tr}}$. Also, the semi-analytic result Eq. (93) is close to the numerical result with $N_{\mathrm{tr}}=7$ around the crossing time. Therefore, we conclude that (i) the crossing condition Eq. (64) can be used for the detection of the crossing of the KL divergence, and (ii) the crossing time can be evaluated by the semi-analytic result Eq. (93).
@@ IV.4 Case (ii) $V(\alpha)=0$
we plot Fig. 9 as the $\beta_{\mathrm{ini}}$ dependence of $a_{m}$ ($m=2,3,4,5$). In contrast to case (i), a peak of $a_{2}$ exists at $\beta_{\mathrm{ini}}=\beta_{\mathrm{ini}}^{*}\approx 1.15>1$. Therefore, if the Mpemba-like effect exists, it should be the inverse Mpemba effect in the heating process.
@@ IV.5 Case (iii) $V(\alpha)>0$
In this case, $a_{2}(\beta_{\mathrm{ini}},\beta)$ decreases with $\beta_{\mathrm{ini}}$ monotonically as shown in Fig. 12(a) while the behaviors of $a_{m}(\beta_{\mathrm{ini}},\beta)$ for $m=2,3,4$ are the similar to the previous cases.
@@ IV.6 Phase diagram of the Mpemba effect
Here, we draw the phase diagram to distinguish the region of the occurrence of the Mpemba effect from the non-Mpemba region under the control of a set of parameters. Figure 13 exhibits a phase diagram on the plane of $(k_{\mathrm{mid}},\alpha)$ by fixing $k_{\mathrm{in}}=k_{\mathrm{out}}=\xi=1$, where the circles and crosses correspond to the occurrence of the Mpemba effect and absence of the Mpemba effect, respectively. We also plot open triangles where $a_{2}$ has a peak, while Eq. (64) is not satisfied. Interestingly, the open triangles are located not only in the vicinity of $V(\alpha)=0$ but also in the region with large $\alpha$. We need to clarify the mechanism for the appearance of such regions.
@@ V.1 Difference between 1D and 2D
Thus, the effective potential mapping onto 1D systems can be written as
@@
has, at least, one zero for $\beta_{\mathrm{ini}}=\beta_{\mathrm{ini}}^{*}$. We demonstrate that $F_{2}(\beta_{\mathrm{ini}})$ has a different sign in the low-temperature and high-temperature limits for the case (i) $V(\alpha)<V(0)=0$, as shown in Appendix D.1. Thus, we conclude the existence $F_{2}(\beta_{\mathrm{ini}}^{*})=0$ at a $\beta_{\mathrm{ini}}=\beta_{\mathrm{ini}}^{*}$. Mathematically, the difference in the Jacobian between 2D and 1D leads to different conclusions, while $r=0$ can be regarded as an effective potential in 1D problem. Interestingly, we can prove the absence of positive $\beta_{\mathrm{ini}}^{*}$ to satisfy $F_{2}(\beta_{\mathrm{ini}}^{*})$ for the case (iii) $V(\alpha)>0$ (see Appendix D.2). When we regard $r=0$ as an effective potential in 1D system, this result is analogous to that in 1D case with a wall on one side [53, 54].
@@ VI Conclusion
By introducing an exactly solvable model, we have obtained exact eigenmodes of the Fokker-Planck equation in a bistable potential. From spectral analysis, we identify the necessary and sufficient conditions for observing the Mpemba effect. Using the slowest and the second slowest eigenmodes of the particle’s motion in a two-dimensional potential, starting from an equilibrium initial condition after a quench into another equilibrium state, we demonstrate the occurrence of a crossing of the KL-divergence, if $V(\alpha)<V(r_{\mathrm{min}})=0$ where $r_{\mathrm{min}}$ is the location of the minimum of $V(r)$ near $r=0$. In contrast to the 1D case, we can observe the Mpemba effect for a particle constrained in a 2D double potential without the introduction of any confined walls because $r=0$ can be regarded as an effective wall. Interestingly, $a_{2}$ has a peak, but neither the Mpemba effect nor the inverse Mpemba effect cannot be observed for $V(\alpha)=0$. For $V(\alpha)>0$, there is no peak for $a_{2}$ and the Mpemba effect cannot be observed.
@@ Appendix A Properties of Tricomi’s $U$-function
The Tricomi (or Tricomi–Confluent) function $U(a,b,z)$ is the second, linearly independent solution of Kummer’s (confluent-hypergeometric) differential equation
@@
which is valid by analytic continuation for non-integer $b$ and extends to integer $b$ by limit. Useful identities and asymptotics are presented as
@@
Appendix B Quantitative Criterion for Crossing of the KL Divergence
@@
Appendix C Contributions of higher-order terms
@@
Appendix D Sign change of $F_{2}(\beta_{\rm ini})$ for the 2D double-well potential
In this appendix, we prove that for the two-dimensional double-well potential defined in Eq. (9), $F_{2}(\beta_{\mathrm{ini}})$ introduced Eq. (75) changes sign between the limits $\beta_{\rm ini}\to 0$ and $\beta_{\rm ini}\to\infty$. Let us rewrite $F_{2}(\beta_{\mathrm{ini}})$ as
@@ Appendix D Sign change of $F_{2}(\beta_{\rm ini})$ for the 2D double-well potential
D.1 (i) The case for $V(\alpha)<V(0)$
Hete, we assume $V(\alpha)<V(0)$ as the case (i).
D.1.1 High-temperature limit $\beta_{\rm ini}\to 0$
@@ D.1.2 Conclusion
By Chebyshev’s integral inequality for similarly ordered functions, for any probability measure (see Appendix D)

### 01 Apr 2026 v3

@@ IV.1.2 $a_{m}$ for $m=2,3,4,5$
Figure 5(a) shows $\beta_{\mathrm{ini}}$ dependence of $a_{2}$ and Fig. 5(b) shows $a_{3}$, $a_{4}$, and $a_{5}$. The results clearly demonstrate the existence of a peak of $a_{2}$ as a function of $\beta_{\mathrm{ini}}$, which is the necessary condition to observe the Mpemba effect. Remarkably, this suggests that the overdamped Langevin system confined in a two-dimensional asymmetric potential exhibits such a peak in $a_{2}$ even without the presence of a hard wall. This result contrasts with that in the one-dimensional asymmetric potential, where the Mpemba effect disappears if the system contains no hard walls [53, 54].
@@ IV.2 Higher-order contributions in the spectrum decomposition
Figure 6(a) shows the initial distribution $P_{\mathrm{ini}}$ given by Eq. (58) for $\beta_{\mathrm{ini}}=0.82$ with various $N_{\mathrm{tr}}(=3,5,7)$. Although the approximation is not accurate for $N_{\mathrm{tr}}=3$, it becomes sufficiently accurate when $N_{\mathrm{tr}}=7$. Figure 6(b) shows the time evolution of the distribution function for $N_{\mathrm{tr}}=7$. For clarity, we plot the difference from the final state $P_{\mathrm{eq}}(r,\beta)$. Since the smallest eigenvalue is $\tilde{\lambda}_{2}=0.20249779968052614$, the distribution function is found to converge to $P_{\mathrm{eq}}(r,\beta)$ for $t\gg 1/\tilde{\lambda}_{2}\approx 5$. Furthermore, as shown in Appendix D, the Mpemba effect persists in terms of the difference in the KL divergence even when higher-order modes are included. Therefore, Eq. (II.3.1) provides a valid description of the dynamics.
@@ IV.3 Condition of the crossing of the KL divergence
Although we have confirmed the existence of a peak of $a_{2}$ as a function of $\beta_{\mathrm{ini}}$, this may not be sufficient to observe the Mpemba effect. To observe the Mpemba effect, we need a crossing condition for the observable. In this subsection, we clarify the sufficient condition to observe the Mpemba effect.
Among many possible observables, we focus on the KL divergence defined as
@@
which is a typical monotonic measure, where $\delta P:=P(\bm{r},t)-P_{\mathrm{eq}}(\bm{r})$. If we are interested in the late stage of relaxation dynamics or the deviation from the initial state and the final state is small, $D^{\mathrm{KL}}(P(t)||P_{\mathrm{eq}})$ can be approximated as
@@
for two different initial conditions A and B. See Appendix C for the derivation of Eq. (63). This quadratic form is valid only when $P(r,t)$ is sufficiently close to equilibrium; in particular, it does not reproduce the exact value at $t=0$ when the initial condition is another equilibrium distribution. Therefore, the condition Eq. (63) should be regarded as a necessary (but not always sufficient) condition for crossing within the validity of the truncated expansion.
@@
We plot the time evolution of the KL divergence $\Delta D(\beta^{\sharp},\beta^{*}):=D^{\mathrm{KL}}(P(\beta^{\sharp},\beta;t)||P_{\mathrm{eq}}(\beta^{\sharp}))-D^{\mathrm{KL}}(P(\beta^{*},\beta;t)||P_{\mathrm{eq}}(\beta))$ during a cooling process starting from $\beta_{\mathrm{ini}}^{\sharp}=0.70$ and $\beta_{\mathrm{ini}}^{*}=0.82$ to $\beta=1$ in Fig. 7 (see the solid line). A clear crossing is observed around $t\approx 1$, which exhibits the Mpemba effect. We now compare the KL divergence, including higher-order terms, by using the expansion of $P(r,t)$ truncated at $N_{\mathrm{tr}}=7$ as the dashed line. The dotted line represents the semi-analytic result of Eq. (97) (see Appendix D). From these results, the two-mode approximation works well, although the crossing time depends on $N_{\mathrm{tr}}$. Also, the semi-analytic result Eq. (97) is close to the numerical result with $N_{\mathrm{tr}}=7$ around the crossing time. Therefore, we conclude that (i) the crossing condition Eq. (63) can be used for the detection of the crossing of the KL divergence, and (ii) the crossing time can be evaluated by the semi-analytic result Eq. (97).
@@ IV.4 Case (ii) $V(\alpha)=0$
we plot Fig. 9 as the $\beta_{\mathrm{ini}}$ dependence of $a_{m}$ ($m=2,3,4,5$). In contrast to case (i), a peak in $a_{2}$ exists at $\beta_{\mathrm{ini}}=\beta_{\mathrm{ini}}^{*}\approx 1.15>1$. Therefore, if the Mpemba-like effect exists, it should be the inverse Mpemba effect in the heating process.
@@ IV.5 Case (iii) $V(\alpha)>0$
In this case, $a_{2}(\beta_{\mathrm{ini}},\beta)$ decreases with $\beta_{\mathrm{ini}}$ monotonically as shown in Fig. 12(a) while the behaviors of $a_{m}(\beta_{\mathrm{ini}},\beta)$ for $m=2,3,4$ are similar to the previous cases.
@@ IV.6 Phase diagram of the Mpemba effect
Here, we draw the phase diagram to distinguish the region of the occurrence of the Mpemba effect from the non-Mpemba region under the control of a set of parameters. Figure 13 exhibits a phase diagram on the plane of $(k_{\mathrm{mid}},\alpha)$ by fixing $k_{\mathrm{in}}=k_{\mathrm{out}}=\xi=1$, where the circles and crosses correspond to the occurrence of the Mpemba effect and absence of the Mpemba effect, respectively. We also plot open triangles where $a_{2}$ has a peak, while Eq. (63) is not satisfied. Interestingly, the open triangles are located not only in the vicinity of $V(\alpha)=0$ but also in the region with large $\alpha$. We note that this open-triangle region remains unchanged even when higher-order terms of the KL divergence are taken into account; specifically, we use the result obtained by setting $N_{\mathrm{tr}}=7$ in Eq. (57) and substituting it into Eq. (59).
@@ V.1 Difference between 1D and 2D
Thus, the effective potential mapping onto one-dimensional systems can be written as
@@
has, at least, one zero for $\beta_{\mathrm{ini}}=\beta_{\mathrm{ini}}^{*}$. We demonstrate that $F_{2}(\beta_{\mathrm{ini}})$ has a different sign in the low-temperature and high-temperature limits for the case (i) $V(\alpha)<V(0)=0$, as shown in Appendix E.1. Thus, we conclude the existence of $F_{2}(\beta_{\mathrm{ini}}^{*})=0$ at a $\beta_{\mathrm{ini}}=\beta_{\mathrm{ini}}^{*}$. Mathematically, the difference in the Jacobian between 2D and 1D leads to different conclusions, while $r=0$ can be regarded as an effective potential in 1D problem. Interestingly, we can prove the absence of positive $\beta_{\mathrm{ini}}^{*}$ to satisfy $F_{2}(\beta_{\mathrm{ini}}^{*})$ for the case (iii) $V(\alpha)>0$ (see Appendix E.2). When we regard $r=0$ as an effective potential in 1D system, this result is analogous to that in 1D case with a wall on one side [53, 54].
@@ VI Conclusion
By introducing an exactly solvable model, we have obtained exact eigenmodes of the Fokker-Planck equation in a bistable potential. From spectral analysis, we identify the necessary and sufficient conditions for observing the Mpemba effect. Using the slowest and the second slowest eigenmodes of the particle’s motion in a two-dimensional potential, starting from an equilibrium initial condition after a quench into another equilibrium state, we demonstrate the occurrence of a crossing of the KL-divergence, if $V(\alpha)<V(r_{\mathrm{min}})=0$ where $r_{\mathrm{min}}$ is the location of the minimum of $V(r)$ near $r=0$. In contrast to the 1D case, we can observe the Mpemba effect for a particle constrained in a 2D double potential without the introduction of any confined walls because $r=0$ can be regarded as an effective wall. Interestingly, $a_{2}$ has a peak, but neither the Mpemba effect nor the inverse Mpemba effect is observed for $V(\alpha)=0$. For $V(\alpha)>0$, there is no peak for $a_{2}$ and the Mpemba effect cannot be observed.
@@ Appendix A Properties of Tricomi’s $U$-function
The Tricomi (or Tricomi-Confluent) function $U(a,b,z)$ is the second, linearly independent solution of Kummer’s (confluent-hypergeometric) differential equation
@@
which is valid by analytic continuation for non-integer $b$ and extends to integer $b$ by limit. Useful identities and asymptotic form are presented as
@@
Appendix B Determination of the eigenmodes
In this appendix, we describe the procedure for determining the eigenmodes. First, we can rewrite Eq. (50) explicitly as
This expression helps one to numerically determine $\lambda$ satisfying Eq. (50).
As explained in the main text, let us focus on the case where the eigenvalues are nontrivial. Let us determine the coefficients $A_{\mathrm{in}}$$B_{\mathrm{out}}$. First, we consider the following set of equations:
$\begin{pmatrix}\Phi^{(1)}_{\rm mid,M}(r_{-})&\Phi^{(1)}_{\rm mid,U}(r_{-})&0\\ \Phi^{(2)}_{\rm mid,M}(r_{-})&\Phi^{(2)}_{\rm mid,U}(r_{-})&0\\ \Phi^{(1)}_{\rm mid,M}(r_{+})&\Phi^{(1)}_{\rm mid,U}(r_{+})&-\Phi^{(1)}_{\rm out,U}(r_{+})\end{pmatrix}\begin{pmatrix}A_{\mathrm{mid}}\\ B_{\mathrm{mid}}\\ B_{\mathrm{out}}\end{pmatrix}=\begin{pmatrix}\Phi^{(1)}_{\rm in,M}(r_{-})\\ \Phi^{(2)}_{\rm in,M}(r_{-})\\ 0\end{pmatrix}A_{\mathrm{in}}.$ (85)
We can easily solve this set of equations as
$\frac{A_{\mathrm{mid}}}{A_{\mathrm{in}}}$ $=\frac{\Phi^{(1)}_{\rm in,M}(r_{-})\Phi^{(2)}_{\rm mid,U}(r_{-})-\Phi^{(2)}_{\rm in,M}(r_{-})\Phi^{(1)}_{\rm mid,U}(r_{-})}{\Phi^{(1)}_{\rm mid,M}(r_{-})\Phi^{(2)}_{\rm mid,U}(r_{-})-\Phi^{(1)}_{\rm mid,U}(r_{-})\Phi^{(2)}_{\rm mid,M}(r_{-})},$ (86)
$\frac{B_{\mathrm{mid}}}{A_{\mathrm{in}}}$ $=\frac{-\Phi^{(1)}_{\rm in,M}(r_{-})\Phi^{(2)}_{\rm mid,M}(r_{-})+\Phi^{(2)}_{\rm in,M}(r_{-})\Phi^{(1)}_{\rm mid,M}(r_{-})}{\Phi^{(1)}_{\rm mid,M}(r_{-})\Phi^{(2)}_{\rm mid,U}(r_{-})-\Phi^{(1)}_{\rm mid,U}(r_{-})\Phi^{(2)}_{\rm mid,M}(r_{-})},$ (87)
$\frac{B_{\mathrm{out}}}{A_{\mathrm{in}}}$ $=\frac{\Phi^{(1)}_{\rm in,M}(r_{-})}{\Phi^{(1)}_{\rm out,U}(r_{+})}\frac{\Phi^{(2)}_{\rm mid,U}(r_{-})\Phi^{(1)}_{\rm mid,M}(r_{+})-\Phi^{(2)}_{\rm mid,M}(r_{-})\Phi^{(1)}_{\rm mid,U}(r_{+})}{\Phi^{(1)}_{\rm mid,M}(r_{-})\Phi^{(2)}_{\rm mid,U}(r_{-})-\Phi^{(1)}_{\rm mid,U}(r_{-})\Phi^{(2)}_{\rm mid,M}(r_{-})}$
$\hskip 10.00002pt+\frac{\Phi^{(2)}_{\rm in,M}(r_{-})}{\Phi^{(1)}_{\rm out,U}(r_{+})}\frac{\Phi^{(1)}_{\rm mid,M}(r_{-})\Phi^{(1)}_{\rm mid,U}(r_{+})-\Phi^{(1)}_{\rm mid,U}(r_{-})\Phi^{(1)}_{\rm mid,M}(r_{+})}{\Phi^{(1)}_{\rm mid,M}(r_{-})\Phi^{(2)}_{\rm mid,U}(r_{-})-\Phi^{(1)}_{\rm mid,U}(r_{-})\Phi^{(2)}_{\rm mid,M}(r_{-})}.$ (88)
Combining Eqs. (86)–(88) with the normalization condition given in Eq. ((56)), one can determine the coefficients $A_{\mathrm{in}}$, $A_{\mathrm{mid}}$, $B_{\mathrm{mid}}$, and $B_{\mathrm{out}}$.
Appendix C Quantitative Criterion for Crossing of the KL Divergence
@@
Appendix D Contributions of higher-order terms
@@
Appendix E Sign change of $F_{2}(\beta_{\rm ini})$ for the 2D double-well potential
In this appendix, we prove that for the two-dimensional double-well potential defined in Eq. (9), $F_{2}(\beta_{\mathrm{ini}})$ introduced in Eq. (74) changes sign between the limits $\beta_{\rm ini}\to 0$ and $\beta_{\rm ini}\to\infty$. Let us rewrite $F_{2}(\beta_{\mathrm{ini}})$ as
@@ Appendix E Sign change of $F_{2}(\beta_{\rm ini})$ for the 2D double-well potential
E.1 (i) The case for $V(\alpha)<V(0)$
Here, we assume $V(\alpha)<V(0)$ as the case (i).
E.1.1 High-temperature limit $\beta_{\rm ini}\to 0$
@@ E.1.2 Conclusion
By Chebyshev’s integral inequality for similarly ordered functions, for any probability measure (see Appendix E)